OpenStudy (anonymous):

Show and explain how replacing one equation by the sum of that equation and a multiple of the other produces a system with the same solutions as the one shown. 8x + 7y = 39 4x – 14y = –68

OpenStudy (anonymous):

hi @justsmile531 <3

OpenStudy (anonymous):

@Lyrae

OpenStudy (anonymous):

multiply the first equation all the way across by \(2\) what do you get?

OpenStudy (anonymous):

16 + 14 ?

OpenStudy (anonymous):

you are missing an \(x\) and a \(y\)

OpenStudy (anonymous):

Oh. so 16x + 14y ?

OpenStudy (anonymous):

better, but you are now missing the equal sign, and the number on the other side don't forget to multiply it by 2 as well

OpenStudy (anonymous):

16y + 14y = 78

OpenStudy (anonymous):

whew now we got it \[16x+14y=79\] is the same line as \[8x+7y=39\] but now we have something better, the system looks like \[16x+14y=79\\ 4x-14y=-68\]

OpenStudy (anonymous):

oops typo there, i meant \[16x+14y=78\\ 4x-14y=-68\]

OpenStudy (anonymous):

now what happens when you add these up in a column?

OpenStudy (anonymous):

Would you add 16x and 4x together? and the same for the rest?

OpenStudy (anonymous):

you would add \(16x+4x\) yes

OpenStudy (anonymous):

that would 20x

OpenStudy (anonymous):

yes it would

OpenStudy (anonymous):

what about the rest?

OpenStudy (anonymous):

Would the 14y + -14y = 0 ?

OpenStudy (anonymous):

yes to that as well so on the left we get just \[20x\] and the \(y\) terms have added up to zero and are gone how about on the right side of the equal sign?

OpenStudy (anonymous):

20x = 10 . Then you divide by 20 on both sides to get x by itself. and you get 2

OpenStudy (anonymous):

slow down there a bit yes you do get \[20x=10\]

OpenStudy (anonymous):

but if you divide by \(20\) you do not get \(2\) \[\frac{20}{10}=2\] but you are computing \[\frac{10}{20}\]

OpenStudy (anonymous):

You get .5 or 1/2

OpenStudy (anonymous):

yes you do

OpenStudy (anonymous):

And now I am done?

OpenStudy (anonymous):

probably need to find \(y\) as well, but that is most of the work done the rest it to replace \(x\) by \(\frac{1}{2}\) in either equation, say \[8x + 7y = 39\] and find \(y\)

OpenStudy (anonymous):

in other words, solve \[8\times \frac{1}{2}+7y=39\] for \(y\)

OpenStudy (anonymous):

y= 5

OpenStudy (anonymous):

bingo

OpenStudy (anonymous):

solution \[x=\frac{1}{2},y=5\] or sometimes written as the pair \[(\frac{1}{2},5)\]

OpenStudy (anonymous):

Wow. Thank you so much. You really helped alot. I appreciate it

OpenStudy (anonymous):

yw, hope it was more or less clear

OpenStudy (anonymous):

You helped a struggling Blonde succeed. Lol

OpenStudy (anonymous):

lol struggling blond was the name of my first punk band

OpenStudy (anonymous):

Nice .

OpenStudy (anonymous):

just kidding though wish it were true

OpenStudy (skullpatrol):

Learning is struggling.

OpenStudy (anonymous):

You got that right.

OpenStudy (skullpatrol):

Great effort!

OpenStudy (anonymous):

Thank you :)